OpenStudy (anonymous):

The wavelength of light in the infrared region is 4.257 x 10^-7 cm. What is the frequency of this light? Need to show work.

OpenStudy (aaronq):

\[\nu=\frac{ c }{ \lambda }\]

OpenStudy (anonymous):

How do you solve it? I've never seen that formula before...

OpenStudy (anonymous):

lights velocity = 3 x 10*8 m/s frequency = v/w = 3 x 10*8/4.257 x 10*-7 = 7.05 x 10*14 Hz